SGI2TIFF(1) SGI2TIFF(1) NAME sgi2tiff−createaTIFFfilefromanSGIimagefile SYNOPSIS sgi2tiff[options]input.rgboutput.tif DESCRIPTION sgi2tiffconvertsafileintheSGIimageformattoTIFF.Bydefault,theTIFFimageiscreated withdatasamplespacked(PlanarConfiguration=1),compressedwiththeLempel-Ziv&Welch algorithm(Compression=5),andwitheachstripnomorethan8kilobytes.Thesecharacteristics canoverriden,orexplicitlyspecifiedwiththeoptionsdescribedbelow. OPTIONS −cSpecifyacompressionschemetousewhenwritingimagedata:−cnonefornocompression,-cpackbitsforthePackBitscompressionalgorithm),-cjpegforthebaseline JPEGcompressionalgorithm,-czipfortheDeflatecompressionalgorithm,and−clzw forLempel-Ziv&Welch(thedefault). −pExplicitlyselecttheplanarconfigurationusedinorganizingdatasamplesintheoutput image:−pcontigforsamplespackedcontiguously,and−pseparateforsamplesstored separately.Bydefaultsamplesarepacked. −rWritedatawithaspecifiednumberofrowsperstrip;bydefaultthenumberofrows/strip isselectedsothateachstripisapproximately8kilobytes. BUGS Doesnotrecordcolormapinformation. SEEALSO tiffinfo(1),tiffcp(1),tiffmedian(1),libtiff(3) October15,1995 1